OpenStudy (anonymous):

Mr. Small, store manager for Jay's Appliance, is having a difficult time placing a selling price on a refrigerator that cost $410. Mr. Small knows his boss would like to have a 45 percent markup based on cost. The selling price should be

OpenStudy (anonymous):

\[410\times 1.45\]

OpenStudy (anonymous):

i get $635.50 with a calculator

OpenStudy (anonymous):

i got 594.5 see ugh

OpenStudy (anonymous):

%45 percent mark up means you want to get %145 of the purchase price. so use a calculator to multiply. or you could find %45 of 410 and then add it to 410 which would give you the same thing.

OpenStudy (anonymous):

%45 of 410 means \[.45\times 410=184.50\]

OpenStudy (anonymous):

oh wait you are right!

OpenStudy (anonymous):

i took 55% instead of 45%! my fault you have it

OpenStudy (anonymous):

1.45 * 410=594.50

OpenStudy (anonymous):

finally just 1

OpenStudy (anonymous):

you are right, i am wrong. punched wrong buttons! good work